the resultant of vectors A and B is 12-N perpendicular to vector A. if vector A also has 12-N magnitude, solve for magnitude and direction of vector B. [Analytical method)
1
Expert's answer
2020-06-01T14:26:31-0400

Solution.

"R=12N;"

"R \perp F_A;"

"F_A=12N;"

"F_B-?;"

"F_B^2=R^2+F_A^2;"

"F_B=\sqrt{R^2+F_A^2};"

"F_B=\sqrt{144N^2+144N^2}=17N;"

"\alpha =45^o+90^o=135^o" -the angle between the vectors A and B;

Answer: "F_B=17N, \alpha = 135^o."